# A note on higher coherence of graphs of groups

### v1

@@ 1. Introduction
For $n\in\mathbb{N}$, a group $G$ is of type $\mathsf{F}_{n}$ (resp. of type $\mathsf{F}_{\infty}$) if there exists a CW-model for the classifying space BGBG with finite $n$-skeleton (resp. of finite type). Every group is of type $\mathsf{F}_{0}$. For a group being of type $\mathsf{F}_{1}$ (resp. of type $\mathsf{F}_{2}$) is equivalent to being finitely generated (resp. finitely presented).
Definition $1.$
@@
Theorem 1.2 was crucial to characterise the coherent right-angled Artin groups. A simplicial complex is flag if every complete subgraph in its 1-skeleton spans a simplex. Associated to a finite flag simplicial complex $L$ is the right-angled Artin group $A_{L}$ that has as generating set the vertices of $L$, subject only to the relations that two generators commute if the corresponding vertices are adjacent in $L$ [Kob22]. A graph is chordal if it does not contain a cycle of length $\geq 4$ as a full subgraph. It is a classical result of Dirac [Dir61] that in chordal graphs, minimal separating sets are cliques. It follows that a flag simplicial complex with chordal 11-skeleton is a “tree of simplices” in the following sense.
@@
    The 11-skeleton of $L$ is a chordal graph;
@@ 2. Inheritance of higher coherence
    If $Q$ is of type $\mathsf{F}_{n}$ and NN is of type $\mathsf{F}_{n}$, then $G$ is of type $\mathsf{F}_{n}$;
@@
    If $G$ is of type $\mathsf{F}_{n}$ and NN is of type $\mathsf{F}_{n-1}$, then $Q$ is of type $\mathsf{F}_{n}$.
@@
    If $Q$ is $(n,m)$-coherent and NN is $(0,m)$-coherent, then $G$ is $(n,m)$-coherent;
@@
    If $G$ is $(n,m)$-coherent and NN is of type $\mathsf{F}_{m-1}$, then $Q$ is $(n,m)$-coherent.
@@
Since NN is $(0,m)$-coherent, the group $N\cap H$ is of type $\mathsf{F}_{m}$ and, in particular, of type $\mathsf{F}_{n-1}$. By Theorem 2.3 ()2.3, the group $p(H)$ is of type $\mathsf{F}_{n}$. Since $Q$ is $(n,m)$-coherent, the group $p(H)$ is of type $\mathsf{F}_{m}$. By Theorem 2.3 ()2.3, it follows that $H$ is of type $\mathsf{F}_{m}$.
@@
In Theorem 1.3 for $n=1$, instead of assuming that the edge groups are (0,m1)(0,m-1)-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with any edge group is of type $\mathsf{F}_{m-1}$. Similarly, in Lemma 2.4 ()2.4, instead of assuming that NN is $(0,m)$-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with NN is of type $\mathsf{F}_{m}$.
@@ 3. Higher coherence of some right-angled Artin groups
We show that $(n,\infty)$-coherence is satisfied by the right-angled Artin groups associated to the flag simplicial complexes in Definition 1.6 ()1.6 and ()1.6 and is preserved under the operations in Definition 1.6 ()1.6 and ()1.6.
@@
(iii) For a gluing $L\coloneqq L_{1}\cup_{L_{0}}L_{2}$ with $L_{1},L_{2}\in\mathcal{T}_{n}$ and $L_{0}\in\mathcal{T}_{1}$ or $L=\emptyset$, the group $A_{L}$ splits as an amalgamated product $A_{L_{1}}\ast_{A_{L_{0}}}A_{L_{2}}$. (Here $A_{\emptyset}$ is the trivial group.) The group $A_{L_{0}}$ is $(1,\infty)$-coherent by Theorem 1.5. By induction on the number of vertices, we may assume that the groups $A_{L_{1}}$ and $A_{L_{2}}$ are $(n,\infty)$-coherent. Then the group $A_{L}$ is $(n,\infty)$-coherent by Theorem 1.3.
(iv) For a cone $CL\coloneqq L\ast\Delta^{0}$ with $L\in\mathcal{T}_{n}$, the group $A_{CL}$ splits as a direct product $A_{L}\times\mathbb{Z}$. By induction on the number of vertices, we may assume that the group $A_{L}$ is $(n,\infty)$-coherent. Then the group $A_{CL}$ is $(n,\infty)$-coherent by Lemma 2.4 ()2.4. ∎
@@
Right-angled Artin groups are so-called graph products of $\mathbb{Z}$. Theorem 1.7 holds more generally for graph products of $(0,\infty)$-coherent groups when Definition 1.6 ()1.6 is modified with respect to their geometric dimension. The $(1,2)$-coherence of graph products was studied in [Var19].
@@
Let $L$ be the flag simplicial complex spanned by the graph in Figure 1. The right-angled Artin group $A_{L}$ is isomorphic to $F_{2}\times F_{2}\times(\mathbb{Z}^{2}\ast\mathbb{Z})$. The group $A_{L}$ is not $(1,2)$-coherent by Theorem 1.5 because $L$ contains a cycle of length 44 as a full subcomplex. The group $A_{L}$ is also not $(2,3)$-coherent by Corollary 3.3 because $L$ is 11-connected and not 22-connected. On the other hand, the group $A_{L}$ is $(4,\infty)$-coherent by Theorem 1.7 because $L$ is 33-dimensional and hence lies in the class $\mathcal{T}_{4}$. We do not know if the group $A_{L}$ is $(3,\infty)$-coherent, although we suspect it to be. Note that $L$ does not lie in the class $\mathcal{T}_{3}$, while every full subcomplex of $L$ that is 22-connected is 33-connected.
$\bullet\bullet\bullet\bullet\bullet\bullet\bullet$
Figure 1.
Remark 3.5.
There is an analogous notion of higher homological coherence [KV23, Fis]. Let $R$ be a ring and let $n\in\mathbb{N}$. A group $G$ is of type $\mathsf{FP}_{n}(R)$ (resp. of type $\mathsf{FP}_{\infty}(R)$) if the trivial RGRG-module $R$ admits a projective resolution that is finitely generated in degrees $\leq n$ (resp. in all degrees).
Let $n,m\in\mathbb{N}\cup\{\infty\}$ with $n<m$. We say that a group $G$ is homologically $(n,m)$-coherent over $R$ if every subgroup of $G$ that is of type $\mathsf{FP}_{n}(R)$ is of type $\mathsf{FP}_{m}(R)$. For $n\geq 2$, homological $(n,m)$-coherence over $\mathbb{Z}$ implies $(n,m)$-coherence because for $n\geq 2$, a group is of type $\mathsf{F}_{n}$ if and only if it is of type $\mathsf{F}_{2}$ and of type $\mathsf{FP}_{n}(\mathbb{Z})$. Inheritance results similar to the ones proved above hold also for higher homological coherence.
Groups of cohomological dimension $n$ over $R$ are homologically $(n,\infty)$-coherent over $R$ [Bro82, Proposition VIII.6.1]. Over the field $\mathbb{Q}$ (and suitably also over other fields), this can be improved if the top-dimensional $\ell^{2}$-Betti number vanishes [Fis, Proposition 2.3]: If $G$ is a locally indicable group with $\operatorname{cd}_{\mathbb{Q}}(G)=n$ and $b^{(2)}_{n}(G)=0$, then $G$ is homologically $(n-1,\infty)$-coherent over $\mathbb{Q}$.
Let $L$ be the flag simplicial complex spanned by the graph in Figure 1. The right-angled Artin group $A_{L}$ satisfies $\operatorname{cd}_{\mathbb{Q}}(A_{L})=4$ and by [DL03]
$b^{(2)}_{4}(A_{L})=\dim_{\mathbb{Q}}H_{3}(L;\mathbb{Q})=0.$
Hence $A_{L}$ is homologically $(3,\infty)$-coherent over $\mathbb{Q}$ (and, in fact, over every field [FHL24, Proposition 3.15][Fis, Proposition 2.3]).
@@ Acknowledgements.
We thank Sam Fisher and Marco Linton for helpful comments. The first author was partially supported by the SFB 1085 Higher Invariants (Universität Regensburg, funded by the DFG).

### v2

@@ 1. Introduction
For $n\in\mathbb{N}$, a group $G$ is of type $\mathsf{F}_{n}$ (resp. of type $\mathsf{F}_{\infty}$) if there exists a CW-model for the classifying space $BG$ with finite $n$-skeleton (resp. of finite type). Every group is of type $\mathsf{F}_{0}$. For a group being of type $\mathsf{F}_{1}$ (resp. of type $\mathsf{F}_{2}$) is equivalent to being finitely generated (resp. finitely presented).
Definition 1.1.
@@
Theorem 1.2 was crucial to characterise the coherent right-angled Artin groups. A simplicial complex is flag if every complete subgraph in its 1-skeleton spans a simplex. Associated to a finite flag simplicial complex $L$ is the right-angled Artin group $A_{L}$ that has as generating set the vertices of $L$, subject only to the relations that two generators commute if the corresponding vertices are adjacent in $L$ [Kob22]. A graph is chordal if it does not contain a cycle of length $\geq 4$ as a full subgraph. It is a classical result of Dirac [Dir61] that in chordal graphs, minimal separating sets are cliques. It follows that a flag simplicial complex with chordal $1$-skeleton is a “tree of simplices” in the following sense.
@@
    The $1$-skeleton of $L$ is a chordal graph;
@@
  3. ()
    All finite flag simplicial complexes $L$ of dimension $n$ with $H_{n}(L;\mathbb{Q})=0$ lie in $\mathcal{T}_{n}$;
@@ 2. Inheritance of higher coherence
    If $Q$ is of type $\mathsf{F}_{n}$ and $N$ is of type $\mathsf{F}_{n}$, then $G$ is of type $\mathsf{F}_{n}$;
@@
    If $G$ is of type $\mathsf{F}_{n}$ and $N$ is of type $\mathsf{F}_{n-1}$, then $Q$ is of type $\mathsf{F}_{n}$.
@@
    If $Q$ is $(n,m)$-coherent and $N$ is $(0,m)$-coherent, then $G$ is $(n,m)$-coherent;
@@
    If $G$ is $(n,m)$-coherent and $N$ is of type $\mathsf{F}_{m-1}$, then $Q$ is $(n,m)$-coherent.
@@
Since $N$ is $(0,m)$-coherent, the group $N\cap H$ is of type $\mathsf{F}_{m}$ and, in particular, of type $\mathsf{F}_{n-1}$. By Theorem 2.3 ()2.3, the group $p(H)$ is of type $\mathsf{F}_{n}$. Since $Q$ is $(n,m)$-coherent, the group $p(H)$ is of type $\mathsf{F}_{m}$. By Theorem 2.3 ()2.3, it follows that $H$ is of type $\mathsf{F}_{m}$.
@@
In Theorem 1.3 for $n=1$, instead of assuming that the edge groups are (0,m1)(0,m-1)-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with any edge group is of type $\mathsf{F}_{m-1}$. Similarly, in Lemma 2.4 ()2.4, instead of assuming that $N$ is $(0,m)$-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with $N$ is of type $\mathsf{F}_{m}$.
3. Higher homological coherence
There is an analogous notion of higher homological coherence [KV23, Fis]. Let $R$ be a ring and let $n\in\mathbb{N}$. A group $G$ is of type $\mathsf{FP}_{n}(R)$ (resp. of type $\mathsf{FP}_{\infty}(R)$) if the trivial $RG$-module $R$ admits a projective resolution that is finitely generated in degrees $\leq n$ (resp. in all degrees).
Definition 3.1.
Let $R$ be a ring and let $n,m\in\mathbb{N}\cup\{\infty\}$ with $n<m$. We say that a group $G$ is homologically $(n,m)$-coherent over $R$ if every subgroup of $G$ that is of type $\mathsf{FP}_{n}(R)$ is of type $\mathsf{FP}_{m}(R)$.
For $n\geq 2$, homological $(n,m)$-coherence over $\mathbb{Z}$ implies $(n,m)$-coherence because for $n\geq 2$, a group is of type $\mathsf{F}_{n}$ if and only if it is of type $\mathsf{F}_{2}$ and of type $\mathsf{FP}_{n}(\mathbb{Z})$. Inheritance results similar to Theorem 1.3 and Lemma 2.4 hold also for higher homological coherence.
Groups of cohomological dimension $n$ over $R$ are homologically $(n,\infty)$-coherent over $R$ [Bro82, Proposition VIII.6.1]. Over the ring $\mathbb{Z}$ (and suitably also over fields), this can be improved if the top-dimensional $\ell^{2}$-Betti number vanishes.
Theorem 3.2 (Fisher, Jaikin-Zapirain–Linton).
Let $n\in\mathbb{N}$ with $n\geq 1$ and let $G$ be a locally indicable group with $\operatorname{cd}_{\mathbb{Z}}(G)=n$. If $b^{(2)}_{n}(G)=0$, then $G$ is homologically $(n-1,\infty)$-coherent over $\mathbb{Z}$.
Proof.
The proof is analogous to that of [Fis, Proposition 2.3], see also [JZL25, Theorem 3.10]. ∎
@@ 4. Higher coherence of some right-angled Artin groups
We show that $(n,\infty)$-coherence is satisfied by the right-angled Artin groups associated to the flag simplicial complexes in Definition 1.6 ()1.6, ()1.6, and ()1.6 and is preserved under the operations in Definition 1.6 ()1.6 and ()1.6.
@@
(iii) If $L$ is of dimension $n$ with $H_{n}(L;\mathbb{Q})=0$, then the group $A_{L}$ satisfies $\operatorname{cd}_{\mathbb{Z}}(A_{L})=n+1$ and $b^{(2)}_{n+1}(A_{L})=\dim_{\mathbb{Q}}H_{n}(L;\mathbb{Q})=0$ by [DL03]. By Theorem 3.2, the group $A_{L}$ is homologically $(n,\infty)$-coherent over $\mathbb{Z}$ and, since $n\geq 2$, hence $(n,\infty)$-coherent.
(iv) For a gluing $L\coloneqq L_{1}\cup_{L_{0}}L_{2}$ with $L_{1},L_{2}\in\mathcal{T}_{n}$ and $L_{0}\in\mathcal{T}_{1}$ or $L_{0}=\emptyset$, the group $A_{L}$ splits as an amalgamated product $A_{L_{1}}\ast_{A_{L_{0}}}A_{L_{2}}$. (Here $A_{\emptyset}$ is the trivial group.) The group $A_{L_{0}}$ is $(1,\infty)$-coherent by Theorem 1.5. By induction on the number of vertices, we may assume that the groups $A_{L_{1}}$ and $A_{L_{2}}$ are $(n,\infty)$-coherent. Then the group $A_{L}$ is $(n,\infty)$-coherent by Theorem 1.3.
(v) For a cone $CL\coloneqq L\ast\Delta^{0}$ with $L\in\mathcal{T}_{n}$, the group $A_{CL}$ splits as a direct product $A_{L}\times\mathbb{Z}$. By induction on the number of vertices, we may assume that the group $A_{L}$ is $(n,\infty)$-coherent. Then the group $A_{CL}$ is $(n,\infty)$-coherent by Lemma 2.4 ()2.4. ∎
@@
Right-angled Artin groups are so-called graph products of $\mathbb{Z}$. Theorem 1.7 holds more generally for graph products of $(0,\infty)$-coherent groups when Definition 1.6 ()1.6 is modified suitably. The $(1,2)$-coherence of graph products was studied in [Var19]. The $\ell^{2}$-Betti numbers and agrarian Betti numbers of graph products were computed in [DO12, FHL24].
@@
Let $L$ be the flag simplicial complex that is the 2-fold suspension of $\Delta^{2}\sqcup\Delta^{0}$. Then $L$ is homotopy equivalent to a 2-dimensional sphere. The right-angled Artin group $A_{L}$ is isomorphic to $F_{2}\times F_{2}\times(\mathbb{Z}^{3}\ast\mathbb{Z})$. The group $A_{L}$ is not $(1,2)$-coherent by Theorem 1.5 because $L$ contains a cycle of length $4$ as a full subcomplex. The group $A_{L}$ is also not $(2,3)$-coherent by Corollary 4.3 because $L$ is $1$-connected and not $2$-connected. On the other hand, the group $A_{L}$ is $(4,\infty)$-coherent by Theorem 1.7 because $L$ is $4$-dimensional with $H_{4}(L;\mathbb{Q})=0$ and hence lies in the class $\mathcal{T}_{4}$. We do not know if the group $A_{L}$ is $(3,\infty)$-coherent. Note that $L$ does not lie in the class $\mathcal{T}_{3}$, while every full subcomplex of $L$ that is $2$-connected is $3$-connected.
@@ Acknowledgements.
We thank Sam Fisher and Marco Linton for very helpful comments. The first author was partially supported by the SFB 1085 Higher Invariants (Universität Regensburg, funded by the DFG).